OpenStudy (anonymous):

(lim h->o) cos((pi/2)+h)/h

OpenStudy (anonymous):

\[\cos\left(\frac{\pi}{2}+h\right)=\cos\frac{\pi}{2}\cos h-\sin\frac{\pi}{2}\sin h=-\sin h\] \[\text{given limit}=-\lim_{h\to0}\frac{\sin h}{h}=-1\]